OpenStudy (anonymous):

HELP Generate the first 5 terms of this sequence f(1)=1, f(2)=2, f(n)=2f(n-2), for n>2 A)1,2,3,4,5 B)1,2,3,5,7 C)1,23,5,8

OpenStudy (gorv):

you r given upto 2

OpenStudy (gorv):

f(n)=2f(n-2) plug in n= 3

OpenStudy (gorv):

n tell what u got ??

OpenStudy (anonymous):

now i got 1,2,3 but what do i do with the n>2

OpenStudy (gorv):

put n=3 in function.... just plug in u will get ans buddy

OpenStudy (gorv):

f(3)=2f(3-2) solve it further i m here to help

OpenStudy (gorv):

@AllanO-the-great

OpenStudy (anonymous):

f(3)=2f is what i got what do i do now @gorv

OpenStudy (gorv):

what is inside f loll f(3)=2*f(1) plug in value of f(1)

OpenStudy (gorv):

f(3)=2*1=2 similarly do for others

OpenStudy (gorv):

put n=4 and calculate and after that put n=5 and calculate

OpenStudy (anonymous):

so its #1 @gorv

OpenStudy (anonymous):

is it #1 Or A) @iGreen

OpenStudy (gorv):

its not in the option